# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
By Theorem D and Lemma 4.8, the slice $S_{\pm}$ is homeomorphic to $\mathrm{Cone}_{0}^{N}(F)$ where $(F,d_{F},\mathfrak{m}_{F})\in\mathsf{RCD}(3-m_{\pm}-2,3-m_{\pm}-1)=\mathsf{RCD}(1-m_{\pm},2-m_{\pm})$, and $(F,d_{F},\mathfrak{m}_{F})$ is a smooth homogeneous space $K_{\pm}/H_{\pm}$. We analyze different cases based on the dimension $m_{\pm}$.
@@
When $m_{\pm}=2$, then $F$ is 00-dimensional, and $Z=\mathrm{Cone}_{0}^{N}(F)$ is an $\mathsf{RCD}(0,1)$-space. By [53], it follows that $Z$ is homeomorphic to one of $\mathbb{R}$, $[0,\infty)$, $[0,1]$, or $\mathbf{S}^{1}$. More over by Lemma 4.8, we have that the tangent cone at $x_{\pm}$ is isometric to $Y\times Z=\mathbb{R}^{2}\times Z$. Since $X$ is non-collapsed we have that the tangent cone at $x_{\pm}$ is isometric to $\mathbb{R}^{3}$. Thus we conclude that $Z$ is isometric to $\mathbb{R}$, and thus the slice $S_{\pm}$ is homeomorphic to $\mathbb{R}$. Moreover, in this case the dimension of $G_{\pm}$ is 00, and thus $G_{\pm}$ is a discrete group acting effectively on $\mathbb{R}$ by homeomorphisms, with $S_{\pm}/G_{\pm}=[0,\varepsilon)$. Thus $G_{\pm}=\mathbb{Z}_{2}$. This implies that $G=T^{2}$, and thus we have that a tubular neighborhood of $G(x_{\pm})$ is homeomorphic to $T^{2}\times_{\mathbb{Z}_{2}}\mathbb{R}$. From the classification in [28] we see that this space is homeomorphic to an Alexandrov space.
@@
When $m_{\pm}=1$, then $F$ is an $\mathsf{RCD}(0,1)$-space, homogeneous, and of dimension $1$. From the characterization of low dimensional $\mathsf{RCD}$-spaces [53] we have that $(F,d_{F},\mathfrak{m}_{F})$ is one of $(\mathbb{R},d_{\mathrm{E}},\mathcal{L}^{1})$, $([0,\infty),d_{\mathrm{E}},h\mathcal{L}^{1})$, $([0,1],d_{\mathrm{E}},h\mathcal{L}^{1})$, $(\mathbf{S}^{1},d_{\mathbf{S}^{1}},h\mathcal{L}^{1})$, where $h$ is a $(0,1)$-convex function. We recall that the due to our non-collapse assumption, the tangent cone at $x_{\pm}$ is isometric to $\mathbb{R}^{3}$. Moreover, due to Lemma 4.4 and Lemma 4.8 we conclude that the slice is homeomorphic to $\mathbb{R}^{2}$, and a cone over $F$. Thus the cases cases when $F$ is homeomorphic to $[0,\infty)$ and $[0,1]$ are excluded. Moreover, the isotropy $G_{\pm}$ acts smoothly by cohomogeneity $1$ (see [29, Corollary E]). Thus we have that $G_{\pm}$ is either $\mathbf{S}^{1}$ when $G=T^{2}$, or $G_{\pm}=\mathrm{SO}(2)$ or $\mathrm{O}(2)$ when $G=\mathrm{SO}(3)$. Comparing to [28], we conclude that the tubular neighborhoods are Alexandrov spaces.
@@
As pointed out in [46, Section 1.6], every compact connected Lie group $G$ has a finite cover of the form $G_{ss}\times T^{k}$, where $G_{ss}$ is semisimple and simply connected. If $G$ acts on $(X,d,m)$ by measure-preserving isometries then every finite cover $\tilde{G}$ of $G$ also acts on $(X,d,m)$ by measure-preserving isometries in the obvious way, but the action of $\tilde{G}$ has non-trivial (but finite) ineffective kernel. Therefore, if we allow for a finite ineffective kernel we can assume $G$ to be of the form $G_{ss}\times T^{k}$.
The classification of semisimple and simply connected compact Lie groups in low dimensions is well known. For the readers convenience we reproduce the part of [46, Table 1.29] that we need below, namely, the list of these groups up to dimension $3$ (there are no semisimple and simply connected compact groups on dimensions $>3$ up to dimension $8$).
@@
$\mathbf{S}^{1}\cong\mathrm{U}(1)\cong\mathrm{SO}(2)$ $1$
$\mathbf{S}^{3}\cong\mathrm{SU}(2)\cong\mathrm{Sp}(1)\cong\mathrm{spin}(3)$ $3$
Let us further observe that although our main topological recognition tool, Theorem D is formulated for effective actions, one obtains the same result in the case of almost effective actions. This follows immediately by observing that from an almost effective action of a Lie group $G$ one gets an effective action of $G/K$ where $K$ is the ineffective kernel of the $G$-action by letting $(gK)x:=gx$ for each $x\in X$. The $G/K$ action then has the same orbits as the $G$-action; consequently the slices of both actions coincide and, since $K$ is finite, the dimensions of $G$ and $G/K$ agree.
In the following theorem we assume as is usual that the action is non-reducible (cf. for example with [45, Definition 2.1.19]) and that the group is acting almost effectively so that we can use the characterization of groups up to cover in the table above. It is worth pointing out that in the list of groups appearing in the corresponding theorem for Alexandrov spaces [28, Theorem C], the identification $\mathrm{SO}(4)\cong\mathbf{S}^{3}\times\mathbf{S}^{3}$ is used. As we are primarily interested in the topological recognition of non-collapsed $\mathsf{RCD}(K,4)$-spaces $X$ we only assume that the action is almost effective, to deal with less possible groups acting on $X$. Note however that, once the classification has been established in this case, one can recover the full classification.
@@
We fix a point $x_{0}$ on a principal orbit and note that $\mathrm{dim}(G(x_{0}))=3$. Now the bound of the dimension estimates (8) grant that $3\leq\mathrm{dim}(G)\leq 6$. As we are allowing for a finite ineffective kernel, the classification of low dimensional compact Lie groups grants that $G$ is, up to finite cover, one of $\mathbf{S}^{3}$, $\mathbf{S}^{3}\times\mathbf{S}^{3}$, $\mathbf{S}^{1}\times\mathbf{S}^{3}$, $T^{2}\times\mathbf{S}^{3}$, $T^{3}\times\mathbf{S}^{3}$, $T^{3}$, $T^{4}$, $T^{5}$ and $T^{6}$. We can immediately rule out the tori $T^{k}$ for $k\geq 4$, as in these cases the principal orbits are homeomorphic to $T^{k}/H$ where $H$ is finite, which is impossible as the space $X$ is $4$-dimensional. Moreover, as we have assumed the action to be non-reducible we can also rule out $T^{2}\times\mathbf{S}^{3}$ by noting that in such a case, we have a $2$-dimensional subgroup $H$ of $\mathbf{S}^{3}\times\{1\}$, which is a contradiction since $\mathbf{S}^{3}$ has no $2$-dimensional subgroups. Finally $T^{3}\times\mathbf{S}^{3}$ is ruled out because the principal isotropy being $3$-dimensional, contains a torus as a proper normal subgroup acting trivially on any point of the space. This contradicts the almost effectiveness of the action. Whence, the first claim is settled.
@@
We consider now $x_{\pm}\in X$ with $x_{\pm}^{\ast}=\pm 1$ and denote by $m_{\pm}=\mathrm{dim}(G(x_{\pm}))$ as in the previous theorem. Observe now that by our hypothesis $N=4$, and that by the Slice Theorem 2.29 we have that small tubular neighborhoods of the orbits $G(x_{\pm})$ are homeomorphic to $G\times_{G_{\pm}}S_{\pm}$, where $S_{\pm}$ is the slice through $x_{\pm}$. Therefore, Theorem D implies that the slice $S_{\pm}$ is homeomorphic to $\mathrm{Cone}_{0}^{N}(F)$ where $(F,d_{F},\mathfrak{m}_{F})\in\mathsf{RCD}(4-m_{\pm}-2,4-m_{\pm}-1)=\mathsf{RCD}(2-m_{\pm},3-m_{\pm})$, and $(F,d_{F},\mathfrak{m}_{F})$ is a smooth homogeneous space $K_{\pm}/H_{\pm}$. We now analyze different cases depending on the value of $m_{\pm}$.
@@
In the case that $m_{\pm}=3$, then $F$ is 00-dimensional. Hence, $Z=\mathrm{Cone}_{0}^{N}(F)$ is an $\mathsf{RCD}(0,1)$-space and, as in the previous theorem, due to [53], is one of $\mathbb{R}$, $[0,\infty)$, $[0,1]$ or $\mathbf{S}^{1}$. Similarly to the proof of the previous theorem, using Lemma 4.8, we get that the tangent cone at $x_{\pm}$ is homeomorphic to $Y\times Z=\mathbb{R}^{3}\times Z$. By the non-collapsed assumption on $X$ we have that the tangent cone at $x_{\pm}$ is homeomorphic to $\mathbb{R}^{4}$. Whence, the slice $S_{\pm}$ is homeomorphic to $\mathbb{R}$. Now again, in this case the dimension of $G_{\pm}$ is zero and therefore $G_{\pm}$ is a discrete group acting effectively on $\mathbb{R}$ by homeomorphisms, with $S_{\pm}/G_{\pm}=[0,\varepsilon)$. This is only possible if $G_{\pm}=\mathbb{Z}_{2}$, in turn forcing $G$ to be either $\mathbf{S}^{3}$ or $T^{3}$. Thus, the small tubular neighborhoods of $G(x_{\pm})$ are equivariantly homeomorphic to either $\mathbf{S}^{3}\times_{\mathbb{Z}_{2}}\mathbb{R}$ or $T^{3}\times_{\mathbb{Z}_{2}}\mathbb{R}$, and it follows from the classification [28] that $X$ is then homeomorphic to an Alexandrov space.
@@
Now we turn to the case in which $m_{\pm}\leq 2$. Under this assumption $2-m_{\pm}\geq 0$ and by [77, Theorem 1], $F$ has a Riemannian metric of positive Ricci curvature if $m_{\pm}\leq 1$ and of non-negative Ricci curvature if $m_{\pm}=2$. Hence, by using [41, Main Theorem], $F$ admits a metric of positive sectional curvature if $m_{\pm}\leq 1$ and of nonnegative sectional curvature if $m_{\pm}=2$. To conclude we only need to observe then that the possible choices of $F$ already appear in the classification [28], and therefore $X$ is homeomorphic to an Alexandrov space. ∎
@@ 7. New Examples of $\mathsf{RCD}$-spaces of cohomogeneity one
Let $F=G/H$ be a homogeneous space of dimension $n\geq 2$ with finite fundamental group. Then there exists a $G$-invariant Riemannian metric $g$ with $\operatorname{Ric}(g)\geq(n-1)>0$ and $\operatorname{diam}(g)\leq\pi$. Moreover, for any $K\geq 0$, the $(K,n)$-cone $\mathrm{Con}^{n}_{K}(F)$ is a non-collapsed $\mathsf{RCD}(Kn,n+1)$ with an action by measure preserving isometries of $G$ making it a cohomogeneity one $\mathsf{RCD}$-space. In the case of $K>0$, the space is homeomorphic to the suspension of $F$ and the orbit space is isometric to $[0,\pi/\sqrt{K}]$ with weighted measure $\sin_{K}(t)^{n}d\mathcal{L}^{1}$ and we have the group diagram $(G,H,G,G)$. In the case when $K=0$, then the space is homeomorphic to the cone over $F$ with the vertex the only fixed point, and the orbit space is isometric to $[0,\infty)$ with weighted measure $t^{n}d\mathcal{L}^{1}$.
@@
By [6, Theorem 1] $F$ admits a $G$-invariant Riemannian metric $\tilde{g}$ with $\operatorname{Ric}(\tilde{g})\geq a>0$. Consider $\lambda>0$ such that $a/(n-1)\geq\lambda^{2}$ and $\pi/\operatorname{diam}(\tilde{g})\geq\lambda$. Then with respect to the Riemannian metric $g=\lambda g$, we have that $(F,d_{g},\mathrm{vol}(g))=(F,\lambda d_{\tilde{g}},\lambda^{n}\mathrm{vol}(\tilde{g}))$ is an $\mathsf{RCD}(a\lambda^{-2},n)$-space. Since by construction $a\lambda^{-2}\geq(n-1)$, then $(F,d_{g},\mathrm{vol}(g))$ is an $\mathsf{RCD}(n-1,n)$-space. By [73, Theorem 2 (i)] it follows that $\operatorname{Ric}(g)\geq(n-1)$. Moreover, by construction we have that $\operatorname{diam}(g)=\lambda\operatorname{diam}(\tilde{g})\leq\pi$. Then by [51, Theorem 1.1] we obtain that for $K\geq 0$ the $(K,n)$-cone $\mathrm{Con}^{n}_{K}(F)$ is $\mathsf{RCD}(Kn,n+1)$. But by construction for $K>0$ the underlying topological space of $\mathrm{Con}^{n}_{K}(F)$ is $F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})$, and thus homeomrphic to the suspension $\mathrm{Susp}(F):=F\times[0,1]/(F\times\{0,\pi/\sqrt{K}\})$. And for $K=0$ the underling space of $\mathrm{Con}_{K}^{n}(F)$ is the topological cone $F\times[0,\infty)/(F\times\{0\})$.
@@
:=$:=\begin{cases}\cos_{K}^{-1}\Big{(}\cos_{K}(s)\cos_{K}(t)+K\sin_{K}(s)\sin_{K}(t)\cos\big{(}d_{F}(x,y)\wedge\pi\big{)}\Big{)}&\mbox{if }K>0,\\ \sqrt{s^{2}+t^{2}-st\cos\big{(}d_{F}(x,y)\wedge\pi\big{)}}&\mbox{if }=0.\end{cases}$
@@
Last we point out that the essential dimension of $\mathrm{Con}^{n}_{K}(F)$ is $n+1$. Thus by [23, Theorem 1.12] and [11, Theorem 1.3] we have that $\mathrm{Con}^{n}_{K}(F)$ is non-collapsed. ∎
Theorem F.

### v3

@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
By ?THM? LABEL:MT:_Geometry_of_the_slice and ?THM? LABEL:L:_infinitesimal_action_cohomogeneity_1, the slice $S_{\pm}$ is homeomorphic to $\mathrm{Cone}_{0}^{N}(F)$ where $(F,d_{F},\mathfrak{m}_{F})\in\mathsf{RCD}(3-m_{\pm}-2,3-m_{\pm}-1)=\mathsf{RCD}(1-m_{\pm},2-m_{\pm})$, and $(F,d_{F},\mathfrak{m}_{F})$ is a smooth homogeneous space $K_{\pm}/H_{\pm}$. We analyze different cases based on the dimension $m_{\pm}$.
@@
When $m_{\pm}=2$, then $F$ is 0-dimensional, and $Z=\mathrm{Cone}_{0}^{N}(F)$ is an $\mathsf{RCD}(0,1)$-space. By [53], it follows that $Z$ is homeomorphic to one of $\mathbb{R}$, $0,\infty)$, $[0,1]$, or $\mathbf{S}^{1}$. More over by ?THM? LABEL:L:_infinitesimal_action_cohomogeneity_1, we have that the tangent space at $x_{\pm}$ is isometric to $Y\times Z=\mathbb{R}^{2}\times Z$. Since $X$ is non-collapsed we have that the tangent space at $x_{\pm}$ is isometric to $\mathbb{R}^{3}$. Thus we conclude that $Z$ is isometric to $\mathbb{R}$, and thus the slice $S_{\pm}$ is homeomorphic to $\mathbb{R}$. Moreover, in this case the dimension of $G_{\pm}$ is 0, and thus $G_{\pm}$ is a discrete group acting effectively on $\mathbb{R}$ by homeomorphisms, with $S_{\pm}/G_{\pm}=[0,\varepsilon)$. Thus $G_{\pm}=\mathbb{Z}_{2}$. This implies that $G=T^{2}$, and thus we have that a tubular neighborhood of $G(x_{\pm})$ is homeomorphic to $T^{2}\times_{\mathbb{Z}_{2}}\mathbb{R}$. From the classification in [26] we see that this space is homeomorphic to an Alexandrov space.
@@
When $m_{\pm}=1$, then $F$ is an $\mathsf{RCD}(0,1)$-space, homogeneous, and of dimension $1$. From the characterization of low dimensional $\mathsf{RCD}$-spaces [53] we have that $(F,d_{F},\mathfrak{m}_{F})$ is one of $(\mathbb{R},d_{\mathrm{E}},\mathcal{L}^{1})$, $([0,\infty),d_{\mathrm{E}},h\mathcal{L}^{1})$, $([0,1],d_{\mathrm{E}},h\mathcal{L}^{1})$, $(\mathbf{S}^{1},d_{\mathbf{S}^{1}},h\mathcal{L}^{1})$, where $h$ is a $(0,1)$-convex function. We recall that the due to our non-collapse assumption, the tangent space at $x_{\pm}$ is isometric to $\mathbb{R}^{3}$. Moreover, due to ?THM? LABEL:L:_Splitting_of_Euclidean_space_in_tangent_cone and ?THM? LABEL:L:_infinitesimal_action_cohomogeneity_1 we conclude that the slice is homeomorphic to $\mathbb{R}^{2}$, and a cone over $F$. Thus the cases cases when $F$ is homeomorphic to $0,\infty)$ and $[0,1]$ are excluded. Moreover, the isotropy $G_{\pm}$ acts smoothly by cohomogeneity $1$ (see [27, Corollary E]). Thus we have that $G_{\pm}$ is either $\mathbf{S}^{1}$ when $G=T^{2}$, or $G_{\pm}=\mathrm{SO}(2)$ or $\mathrm{O}(2)$ when $G=\mathrm{SO}(3)$. Comparing to [26], we conclude that the tubular neighborhoods are Alexandrov spaces.
@@
As pointed out in [45, Section 1.6], every compact connected Lie group $G$ has a finite cover of the form $G_{ss}\times T^{k}$, where $G_{ss}$ is semisimple and simply connected. If $G$ acts on $(X,d,m)$ by measure-preserving isometries then every finite cover $\tilde{G}$ of $G$ also acts on $(X,d,m)$ by measure-preserving isometries in the obvious way, but the action of $\tilde{G}$ has non-trivial (but finite) ineffective kernel. Therefore, if we allow for a finite ineffective kernel we can assume $G$ to be of the form $G_{ss}\times T^{k}$.
The classification of semisimple and simply connected compact Lie groups in low dimensions is well known. For the readers convenience we reproduce the part of [45, Table 1.29] that we need below, namely, the list of these groups up to dimension $3$ (there are no semisimple and simply connected compact groups on dimensions >3>3 up to dimension $8$).
@@
$\mathbf{S}^{1}\cong\mathrm{U}(1)\cong\mathrm{SO}(2)$ 11
$\mathbf{S}^{3}\cong\mathrm{SU}(2)\cong\mathrm{Sp}(1)\cong\mathrm{spin}(3)$ 33
Let us further observe that although our main topological recognition tool,?THM? LABEL:MT:_Geometry_of_the_slice is formulated for effective actions, one obtains the same result in the case of almost effective actions. This follows immediately by observing that from an almost effective action of a Lie group $G$ one gets an effective action of $G/K$ where $K$ is the ineffective kernel of the $G$-action by letting $(gK)x:=gx$ for each $x\in X$. The $G/K$ action then has the same orbits as the $G$-action; consequently the slices of both actions coincide and, since $K$ is finite, the dimensions of $G$ and $G/K$ agree.
In the following theorem we assume as is usual that the action is non-reducible (cf. for example with [44, Definition 2.1.19]) and that the group is acting almost effectively so that we can use the characterization of groups up to cover in the table above. It is worth pointing out that in the list of groups appearing in the corresponding theorem for Alexandrov spaces [26, Theorem C], the identification $\mathrm{SO}(4)\cong\mathbf{S}^{3}\times\mathbf{S}^{3}$ is used. As we are primarily interested in the topological recognition of non-collapsed $\mathsf{RCD}(K,4)$-spaces $X$ we only assume that the action is almost effective, to deal with less possible groups acting on $X$. Note however that, once the classification has been established in this case, one can recover the full classification.
@@
We fix a point $x_{0}$ on a principal orbit and note that $\mathrm{dim}(G(x_{0}))=3$. Now the bound of the dimension (10) grants that $3\leq\mathrm{dim}(G)\leq 6$. As we are allowing for a finite ineffective kernel, the classification of low dimensional compact Lie groups grants that $G$ is, up to finite cover, one of $\mathbf{S}^{3}$, $\mathbf{S}^{3}\times\mathbf{S}^{3}$, $\mathbf{S}^{1}\times\mathbf{S}^{3}$, $T^{2}\times\mathbf{S}^{3}$, $T^{3}\times\mathbf{S}^{3}$, $T^{3}$, $T^{4}$, $T^{5}$ and $T^{6}$. We can immediately rule out the tori $T^{k}$ for $k\geq 4$, as in these cases the principal orbits are homeomorphic to $T^{k}/H$ where $H$ is finite, which is impossible as the space $X$ is $4$-dimensional. Moreover, as we have assumed the action to be non-reducible we can also rule out $T^{2}\times\mathbf{S}^{3}$ by noting that in such a case, we have a $2$-dimensional subgroup $H$ of $\mathbf{S}^{3}\times\{1\}$, which is a contradiction since $\mathbf{S}^{3}$ has no $2$-dimensional subgroups. Finally $T^{3}\times\mathbf{S}^{3}$ is ruled out because the principal isotropy being $3$-dimensional, contains a torus as a proper normal subgroup acting trivially on any point of the space. This contradicts the almost effectiveness of the action. Whence, the first claim is settled.
@@
We consider now $x_{\pm}\in X$ with $x_{\pm}^{\ast}=\pm 1$ and denote by $m_{\pm}=\mathrm{dim}(G(x_{\pm}))$ as in the previous theorem. Observe now that by our hypothesis $N=4$, and that by the Slice ?THM? LABEL:T:Slice-Theorem we have that small tubular neighborhoods of the orbits $G(x_{\pm})$ are homeomorphic to $G\times_{G_{\pm}}S_{\pm}$, where $S_{\pm}$ is the slice through $x_{\pm}$. Therefore, ?THM? LABEL:MT:_Geometry_of_the_slice implies that the slice $S_{\pm}$ is homeomorphic to $\mathrm{Cone}_{0}^{N}(F)$ where $(F,d_{F},\mathfrak{m}_{F})\in\mathsf{RCD}(4-m_{\pm}-2,4-m_{\pm}-1)=\mathsf{RCD}(2-m_{\pm},3-m_{\pm})$, and $(F,d_{F},\mathfrak{m}_{F})$ is a smooth homogeneous space $K_{\pm}/H_{\pm}$. We now analyze different cases depending on the value of $m_{\pm}$.
@@
In the case that $m_{\pm}=3$, then $F$ is 0-dimensional. Hence, $Z=\mathrm{Cone}_{0}^{N}(F)$ is an $\mathsf{RCD}(0,1)$-space and, as in the previous theorem, due to [53], is one of $\mathbb{R}$, $0,\infty)$, $[0,1]$ or $\mathbf{S}^{1}$. Similarly to the proof of the previous theorem, using ?THM? LABEL:L:_infinitesimal_action_cohomogeneity_1, we get that the tangent space at $x_{\pm}$ is homeomorphic to $Y\times Z=\mathbb{R}^{3}\times Z$. By the non-collapsed assumption on $X$ we have that the tangent space at $x_{\pm}$ is homeomorphic to $\mathbb{R}^{4}$. Whence, the slice $S_{\pm}$ is homeomorphic to $\mathbb{R}$. Now again, in this case the dimension of $G_{\pm}$ is zero and therefore $G_{\pm}$ is a discrete group acting effectively on $\mathbb{R}$ by homeomorphisms, with $S_{\pm}/G_{\pm}=[0,\varepsilon)$. This is only possible if $G_{\pm}=\mathbb{Z}_{2}$, in turn forcing $G$ to be either $\mathbf{S}^{3}$ or $T^{3}$. Thus, the small tubular neighborhoods of $G(x_{\pm})$ are equivariantly homeomorphic to either $\mathbf{S}^{3}\times_{\mathbb{Z}_{2}}\mathbb{R}$ or $T^{3}\times_{\mathbb{Z}_{2}}\mathbb{R}$, and it follows from the classification [26] that $X$ is then homeomorphic to an Alexandrov space.
@@
Now we turn to the case in which $m_{\pm}\leq 2$. Under this assumption $2-m_{\pm}\geq 0$ and by [80, Theorem 1], $F$ has a Riemannian metric of positive Ricci curvature if $m_{\pm}\leq 1$ and of non-negative Ricci curvature if $m_{\pm}=2$. Hence, by using [39, Main Theorem], $F$ admits a metric of positive sectional curvature if $m_{\pm}\leq 1$ and of nonnegative sectional curvature if $m_{\pm}=2$. To conclude we only need to observe then that the possible choices of $F$ already appear in the classification [26], and therefore $X$ is homeomorphic to an Alexandrov space. ∎
@@ 7. New Examples of $\mathsf{RCD}$-spaces of cohomogeneity one
Let $F=G/H$ be a homogeneous space of dimension $n\geq 2$ with finite fundamental group. Then there exists a $G$-invariant Riemannian metric $g$ with $\operatorname{Ric}(g)\geq(n-1)>0$ and $\operatorname{diam}(g)\leq\pi$. Moreover, for any $K\geq 0$, the $(K,n)$-cone $\mathrm{Con}^{n}_{K}(F)$ is a non-collapsed $\mathsf{RCD}(Kn,n+1)$ with an action by measure preserving isometries of $G$ making it a cohomogeneity one $\mathsf{RCD}$-space. In the case of $K>0$, the space is homeomorphic to the suspension of $F$ and the orbit space is isometric to $0,\pi/\sqrt{K}]$ with weighted measure $\sin_{K}(t)^{n}d\mathcal{L}^{1}$ and we have the group diagram $(G,H,G,G)$. In the case when $K=0$, then the space is homeomorphic to the cone over $F$ with the vertex the only fixed point, and the orbit space is isometric to $0,\infty)$ with weighted measure $t^{n}d\mathcal{L}^{1}$.
@@
By [9, Theorem 1] $F$ admits a $G$-invariant Riemannian metric $\tilde{g}$ with $\operatorname{Ric}(\tilde{g})\geq a>0$. Consider $\lambda>0$ such that $a/(n-1)\geq\lambda^{2}$ and $\pi/\operatorname{diam}(\tilde{g})\geq\lambda$. Then with respect to the Riemannian metric $g=\lambda g$, we have that $)(F,d_{g},\mathrm{vol}(g))=(F,\lambda d_{\tilde{g}},\lambda^{n}\mathrm{vol}(\tilde{g}$)) is an $\mathsf{RCD}(a\lambda^{-2},n)$-space. Since by construction $a\lambda^{-2}\geq(n-1)$, then $(F,d_{g},\mathrm{vol}(g))$ is an $\mathsf{RCD}(n-1,n)$-space. By [77, Theorem 2 (i)] it follows that $\operatorname{Ric}(g)\geq(n-1)$. Moreover, by construction we have that $\operatorname{diam}(g)=\lambda\operatorname{diam}(\tilde{g})\leq\pi$. Then by [50, Theorem 1.1] we obtain that for $K\geq 0$ the $(K,n)$-cone $\mathrm{Con}^{n}_{K}(F)$ is $\mathsf{RCD}(Kn,n+1)$. But by construction for $K>0$ the underlying topological space of $\mathrm{Con}^{n}_{K}(F)$ is $F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})$, and thus homeomorphic to the suspension $\mathrm{Susp}(F):=F\times[0,1]/(F\times\{0,\pi/\sqrt{K}\})$. And for $K=0$ the underling space of $\mathrm{Con}_{K}^{n}(F)$ is the topological cone $F\times[0,\infty)/(F\times\{0\})$.
@@
$:=\begin{cases}\cos_{K}^{-1}\Big(\cos_{K}(s)\cos_{K}(t)+K\sin_{K}(s)\sin_{K}(t)\cos\big(d_{F}(x,y)\wedge\pi\big)\Big)&\mbox{if }K>0,\\ \sqrt{s^{2}+t^{2}-st\cos\big(d_{F}(x,y)\wedge\pi\big)}&\mbox{if }K=0.\end{cases}$
@@
Last we point out that the essential dimension of $\mathrm{Con}^{n}_{K}(F)$ is $n+1$. Thus by [22, Theorem 1.12] and [12, Theorem 1.3] we have that $\mathrm{Con}^{n}_{K}(F)$ is non-collapsed. ∎
Theorem LABEL:MT:_Grassmannians_as_examples.